# The near horizon dynamics of three-dimensional Einstein gravity

### v1

@@ 5 Euclidean theory
such that $f(\tau+\beta)=f(\tau)+\beta$. As an example the form of the gauge field $A_{\tau}\equiv-iA_{u}$ using the analytic continuation (5) is
$A_{\tau}$ $=aL_{0}-\frac{i}{2}TL_{-1}-iM_{1}-\frac{i}{\ell}\,.$ (5.65)
We will use the Euclidean $A_{\tau}$ for holonomy considerations in the following section for defining a black hole in the theory through its properties, such as regularity at the horizon in Euclidean spacetime.
$\tau$$x$
@@ 5.1 Holonomy condition
Our background has a horizon; thus, we should impose horizon regularity conditions in Euclidean signature. Translating this condition for the Euclidean gauge connection $A_{\tau}$ for the boosted Rindler AdS3 (5.65) or the boosted Rindler flat ($\ell\to\infty$), amounts to imposing the triviality of the holonomy of the Euclidean connection along contractible cycles Gutperle:2011kf ; Ammon:2011nk ; Kraus:2011ds which is $\tau$ in our case. This necessitates the holonomy as a group element to belong to the center of the group Bunster:2014mua ;
@@
where ${\mathbb{Z}}_{2}$ is the centre of Spin$(2,2)\simeq\text{SL}(2,\mathbb{R})\times\text{SL}(2,\mathbb{R})$. Thus both elements $\{I,-I\}\in{\mathbb{Z}}_{2}$ are trivial elements in the actual SO(2,2)22(2,2)( 2 , 2 ) group. This amounts to consider the holonomy Hol[$A_{\tau}$] in (5.66) to be trivial in SO(2,2)22(2,2)( 2 , 2 ) but nontrivial in Spin(2,2)22(2,2)( 2 , 2 ) as it corresponds to minus identity. In other words, we require the configuration (2.8) to be regular (single-valued) as a gauge connection in SO(2,2)22(2,2)( 2 , 2 ) group but not in Spin(2,2)22(2,2)( 2 , 2 ) Castro:2011iw ; Cotler:2018zff ; Cotler:2020ugk .
@@
Accordingly, the holonomy is a center element in both ISO$(2,1)$ and ISL$(2,\mathbb{R})$ but by requiring the connection to be regular only in the former and not in the latter; Hol$[A_{\tau}]=-I$. In fact for the Rindler space which is the vacuum of our theory with $(a_{0},\eta_{0})=(\frac{2\pi}{\beta},0)$ the holonomy of the Euclidean connection is $-I$ as we notice below.
@@
In the finite $\ell$ case corresponding to the boosted Rindler AdS3 configuration (4.45) evaluating the center independent of any chosen representation is hard. Since the algebra is semi-simple, we choose a representation and evaluate the holonomy accordingly. The fundamental representation of the so(2,2)22(2,2)( 2 , 2 ) algebra (2.7) consists of six $4\times 4$ matrices as [^6] We notice that the non-degenerate metric using the ordinary trace in this representation does not correspond to the bilinear form below (2.7), in fact it reproduces the exotic bilinear form Witten:1988hc . However, since we do not use any trace in calculating the holonomy, it is safe to use this representation for our purpose.
@@ 5.2 Flat Rindler boundary action
$I^{\text{\tiny bRflat}}[h,\alpha]=\frac{k}{8\pi}\int\operatorname{d}\!\tau\operatorname{d}\!x\Big{(}\frac{2\pi iN}{\beta}(\dot{h}\alpha^{\prime}+\dot{\alpha}h^{\prime})+\dot{\alpha}\frac{\dot{h}^{\prime}}{\dot{h}}+\alpha^{\prime}\frac{\ddot{h}}{\dot{h}}+\frac{2\pi iN}{\beta}\dot{h}+\frac{\ddot{h}}{\dot{h}}\Big{)}\,.$ (5.82)
@@
$I^{\text{\tiny bRflat}}[h,\alpha]=\frac{k}{8\pi}\int\operatorname{d}\!\tau\operatorname{d}\!x\Big{(}-2\alpha\big{(}\frac{2\pi iN}{\beta}\dot{h}+\frac{\ddot{h}}{\dot{h}}\big{)}^{\prime}+\frac{2\pi iN}{\beta}\dot{h}+\frac{\ddot{h}}{\dot{h}}\Big{)}\,.$ (5.83)
@@ 5.3 Rindler AdS3 boundary action
$I^{\text{\tiny bRAdS}}[h,\alpha]=I^{\text{\tiny bRflat}}[h,\alpha]+\frac{k}{4\pi}\int\operatorname{d}\!\tau\operatorname{d}\!x\Big{[}\pm\frac{4\pi^{2}\ell}{\beta^{2}}N_{3}\dot{h}h^{\prime}\pm\frac{\pi i\ell N_{4}}{2\beta}\left(\frac{\ddot{h}\,h^{\prime}}{\dot{h}}+\dot{h}^{\prime}\right)-\frac{\dot{\alpha}\alpha^{\prime}}{\ell}-\frac{\dot{\alpha}}{\ell}\Big{]}\,.$ (5.87)
@@ 6 Discussion
We worked out the boundary effective action governing the dynamics of 2D allowed diffeomorphisms corresponding to the field theory of the boundary gravitons on the 3D Rindler spacetime. Since our theory always involves a horizon, in the Euclidean theory we imposed the regularity condition at the horizon via imposing triviality of the holonomy conditions. This fixed the zero modes corresponding to the exceptional coadjoint orbits of the asymptotic symmetry group which is the warped Virasoro group.
@@ Acknowledgment
NA would like to thank Pavel Putrov, Monica Guica and Cumrun Vafa for discussions on general aspects of holography and the Hamiltonian reduction in 3D gravity during her participation in the Spring School on Superstring Theory and Related Topics 2023, the New Pathways in Exploration of Quantum Field Theory 2023, and Quantum Gravity beyond Supersymmetry III and String-Math 2024 at ICTP. Her appreciation also goes to the Abdus Salam International Centre for Theoretical Physics (ICTP) for providing financial support throughout her stay. This research was supported by the Iran National Science Foundation (INSF), project No. 4000132.

### v4

@@ 5 Euclidean theory
and that $f(\tau+\beta)=f(\tau)+\beta$. As an example the form of the gauge field $A_{\tau}\equiv-iA_{u}$ using the analytic continuation (5) is
$A_{\tau}$ $=aL_{0}-\frac{i}{2}TL_{-1}-iM_{1}-\frac{i}{\ell}L_{1}\,.$ (5.65)
We will use the Euclidean $A_{\tau}$ for holonomy considerations in the following section for defining a black hole in the theory through its properties, such as regularity at the horizon in Euclidean spacetime;
@@ 5.1 Holonomy condition
Our background has a horizon; thus, we should impose horizon regularity conditions in Euclidean signature. Translating this condition for the Euclidean gauge connection $A_{\tau}$ for the boosted Rindler AdS3 (5.65) or the boosted Rindler flat ($\ell\to\infty$), amounts to imposing the triviality of the holonomy of the Euclidean connection along contractible cycles Gutperle:2011kf ; Ammon:2011nk ; Kraus:2011ds which is $\tau$ in our case. This necessitates the holonomy as a group element belonging to the center of the group Bunster:2014mua ;
@@
where ${\mathbb{Z}}_{2}$ is the centre of Spin$(2,2)\simeq\text{SL}(2,\mathbb{R})\times\text{SL}(2,\mathbb{R})$. Thus both elements $\{\mathds{1},-\mathds{1}\}\in{\mathbb{Z}}_{2}$ are trivial elements in the actual SO(2,2)22(2,2)( 2 , 2 ) group. This amounts to consider the holonomy Hol[$A_{\tau}$] in (5.66) to be trivial in SO(2,2)22(2,2)( 2 , 2 ) but nontrivial in Spin(2,2)22(2,2)( 2 , 2 ) as it corresponds to minus identity. In other words, we require the configuration (2.8) to be regular (single-valued) as a gauge connection in SO(2,2)22(2,2)( 2 , 2 ) group but not in Spin(2,2)22(2,2)( 2 , 2 ) Castro:2011iw ; Cotler:2018zff ; Cotler:2020ugk .
@@
Accordingly, the holonomy is a center element in both ISO$(2,1)$ and ISL$(2,\mathbb{R})$ but by requiring the connection to be regular only in the former and not in the latter; Hol$[A_{\tau}]=-\mathds{1}$. In fact for the Rindler space which is the vacuum of our theory with $(a_{0},\eta_{0})=(\frac{2\pi}{\beta},0)$ the holonomy of the Euclidean connection is $-\mathds{1}$ as we notice below.
@@
In the finite $\ell$ case corresponding to the boosted Rindler AdS3 configuration (4.45), evaluating the center of the group, independent of any chosen representation is hard. Since the algebra is semi-simple, we choose a representation and evaluate the holonomy accordingly. The fundamental representation of the so(2,2)22(2,2)( 2 , 2 ) algebra (2.7) consists of six $4\times 4$ matrices as [^6] We notice that the non-degenerate metric using the ordinary trace in this representation does not correspond to the bilinear form below (2.7), in fact it reproduces the exotic bilinear form Witten:1988hc . However, since we do not use any trace in calculating the holonomy, it is safe to use this representation for our purpose.
@@ 5.2 Flat Rindler boundary action
$I^{\text{\tiny Rflat}}[h,\alpha]=\frac{k}{4\pi}\int\operatorname{d}\!\tau\operatorname{d}\!x\Big{(}\frac{2\pi iN}{\beta}(\dot{h}\alpha^{\prime}+\dot{\alpha}h^{\prime})+\dot{\alpha}\frac{\dot{h}^{\prime}}{\dot{h}}+\alpha^{\prime}\frac{\ddot{h}}{\dot{h}}+\frac{4\pi iN}{\beta}\dot{h}+2\frac{\ddot{h}}{\dot{h}}\Big{)}\,,$ (5.82)
@@
$I^{\text{\tiny Rflat}}[h,\alpha]=\frac{k}{2\pi}\int\operatorname{d}\!\tau\operatorname{d}\!x\Big{(}-\alpha\big{(}\frac{2\pi iN}{\beta}\dot{h}+\frac{\ddot{h}}{\dot{h}}\big{)}^{\prime}+\frac{2\pi iN}{\beta}\dot{h}+\frac{\ddot{h}}{\dot{h}}\Big{)}\,.$ (5.83)
@@ 5.3 Rindler AdS3 boundary action
$I^{\text{\tiny RAdS}}[h,\alpha]=I^{\text{\tiny Rflat}}[h,\alpha]+\frac{k}{2\pi}\int\operatorname{d}\!\tau\operatorname{d}\!x\Big{[}\pm\frac{4\pi^{2}\ell}{\beta^{2}}N_{3}\dot{h}h^{\prime}\pm\frac{\pi i\ell N_{4}}{2\beta}\left(\frac{\ddot{h}\,h^{\prime}}{\dot{h}}+\dot{h}^{\prime}\right)-\frac{\dot{\alpha}\alpha^{\prime}}{\ell}-\frac{2\dot{\alpha}}{\ell}\Big{]}\,,$ (5.87)
@@ 6 Discussion
We worked out the boundary effective action governing the dynamics of 2D allowed diffeomorphisms corresponding to the field theory of the boundary gravitons on the 3D Rindler spacetime. Since our theory always involves a horizon, in the Euclidean theory we imposed the regularity condition at the horizon via imposing triviality of the holonomy conditions. This fixed the zero modes of state-dependent functions $(a,T)$ corresponding to the exceptional coadjoint orbits of the asymptotic symmetry group, the warped Virasoro group.
Entropy.
The entropy density $s=S/\Delta$ with $\Delta=\int_{-\infty}^{\infty}\operatorname{d}\!x$ in the canonical ensemble can be calculated using the on-shell Euclidean action, which in this case using (5.83) and introducing $I^{\text{\tiny E}}\equiv iI_{\text{\tiny Wick-rot.}}$ leads to [^7] This convention guarantees that the real contour for the Lagrange multiplier $\alpha$ leads to a well-defined Euclidean path integral.
$S=-I^{\text{\tiny E}}_{\text{\tiny on-shell}}=\frac{-ik}{2\pi}\int_{-\infty}^{\infty}\operatorname{d}\!x\int_{0}^{\beta}\operatorname{d}\!\tau\Big{(}\frac{2\pi iN}{\beta}\dot{h}+\frac{\ddot{h}}{\dot{h}}\Big{)}=\frac{-ik\Delta}{2\pi}\left(2\pi iN\right)=\frac{N\Delta}{4G}\,,$ (6.88)
where in the last identity we used that $k=1/(4G)$. We note that $N$ is an odd number with $N=1$ for the Rindler background leading to $s=1/(4G)$. For Rindler thermodynamics see Laflamme:1987ec and appendix B of Afshar:2015wjm .
@@ Acknowledgment
We thank the anonymous referee for his/her valuable comments. We thank Daniel Grumiller for his comments on the draft. NA would like to thank Monica Guica, Pavel Putrov and Cumrun Vafa for discussions on general aspects of holography and the Hamiltonian reduction in 3D gravity during her participation in the Spring School on Superstring Theory and Related Topics 2023, the New Pathways in Exploration of Quantum Field Theory 2023, and Quantum Gravity beyond Supersymmetry III and String-Math 2024 at ICTP. She also appreciates the Abdus Salam International Centre for Theoretical Physics (ICTP) for providing financial support throughout her stay. This research was supported by the Iran National Science Foundation (INSF), project No. 4000132.